# Milnor number of invariant singularity: generalization of Chulkov’s inequality

### v1

@@ Abstract.
In [3] S.P. Chulkov has proven that if the Taylor series of $f$ contains no monomials of degree $\leq 2$ then $\mu(f)\geq p-1$ for any function germ invariant with respect to a nontrivial action of $\mathbb{Z}_{p}$ for prime pp. Here we present a generalization of this inequality for actions of arbitrary abelian groups.
@@ 1. Introduction
Theorem $1.$
Consider a germ of function $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ with an isolated singularity at the origin and a compact Lie group $G$ of analytic diffeomorphisms $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{n},0)$ such that $f(g(x))=f(x)\;\forall\;g\in G$. If the Taylor series of $f$ contains no monomials of degree $\leq 2$ then $G$ is discrete.
@@ 2. Definitions and auxiliary statements
A function $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ is called equivariant with respect to group action $\tau$ on $(\mathbb{C}^{n},0)$ if there is a character $\lambda$ of $G$ such that $f(\tau(g)\circ x)=\lambda(g)f(x)\;\forall\;x$. $\lambda$ is called the weight of $f$. Invariant functions are equivariant with weight 1, of course.
@@
If $f$ is invariant with respect to an action of a group $G$, then $G$ also acts on $Q_{f}$. The representation of $G$ on $Q_{f}$ is called the equvariant Milnor number of $f$ and denoted $\mu_{G}(f)$. $\nu(f)$ is the multiplicity of the trivial representation in $\mu_{G}(f)$, i.e. the dimension of the space of invariant elements of $Q_{f}$. This subspace of invariant elements is denoted $Q^{G}_{f}$.
@@
The set $C$ of critical points of any invariant deformation $f_{\lambda}$ is $G$-invariant. The action of $G$ on $(\mathbb{C}^{n},0)$ induces a representation $\theta$ of $G$ on the space of $\mathbb{C}$-valued functions on $C$.
Theorem $2.$
@@
Consider the action $\tau:g\circ(x_{1},\dots,x_{n})\longrightarrow\\ \longrightarrow(x_{1},\dots,x_{k},\chi_{k+1}(g)x_{k+1},\dots,\chi_{n}(g)x_{n}),\chi_{k+1},\dots,\chi_{n}\neq 1$ and an invariant function $f$ with a zero 2jet2-jet and an isolated critical point and the origin.
@@ 3. Proof of main theorem
For a group action $\tau:g\circ(x_{1},\dots,x_{n})\longrightarrow(\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n})$ we will denote $\tau_{\mathbb{R}}$ the action on $\mathbb{C}^{2n}$ defined as $g\circ(x_{1},\dots,x_{n},y_{1},\dots,y_{n})\longrightarrow\\ (\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n},\bar{\chi}_{1}(g)y_{1},\dots,\bar{\chi}_{n}(g)y_{n})$. $\tau_{\mathbb{R}}$ is always a real action with the invariant quadratic form $\Sigma_{i}x_{i}y_{i}$. If $\tau$ has no fixed points outside of origin then neither does $\tau_{\mathbb{R}}$.
For a function $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0) ff$$\oplus f$ is the function $f(x_{1},\dots,x_{n})+f(y_{1},\dots,y_{n})$. It is easy to see that if $f$ has an isolated singularity at the origin then so does $f\oplus f$, and $\mu(f\oplus f)=(\mu(f))^{2}$. Furthermore, if $f$ is invariant with respect to an action $\tau$ then $f\oplus f$ is invariant with respect to the action $\tau_{\mathbb{R}}$.
@@
Now we can conclude the proof of the main theorem. As $\tau_{\mathbb{R}}$ is a real representation, we can use Roberts’ inequality for the Milnor number of $f\oplus f$: $\mu(f\oplus f)\geq(\nu(f\oplus f)-1)l(\tau_{\mathbb{R}})+1$.
As we now know $\mu(f\oplus f)=\mu^{2}(f)$, $\nu(f\oplus f)\geq\mu(f)$ and, obviously, $l(\tau_{\mathbb{R}})=l(\tau)$. Substituting all of these into Roberts’ inequalit gives us the inequality $\mu^{2}(f)\geq(\mu(f)-1)l(\tau)+1$.

### v3

# Milnor number of an invariant singularity: generalization of Chulkov’s inequality

@@ Abstract.
S.P. Chulkov has proven that the Milnor number of a function germ $f$ with zero 2-jet invariant with respect to a nontrivial action of $\mathbb{Z}_{p}$ is at least p1p-1 for prime $p$. In this paper we prove a generalization of this inequality for actions of arbitrary abelian groups using M. Roberts’ results on invariant morsifications for invariants of real group actions.
@@ 1. Introduction
Theorem 1.1.
Consider a germ of function $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ with an isolated singularity at the origin and a compact Lie group $G$ of analytic diffeomorphisms $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{n},0)$ such that $f(g(x))=f(x)\;\forall\;g\in G\;\forall\;x\in\mathbb{C}^{n}$. If the Taylor series of $f$ contains no monomials of degree $\leq 2$ then $G$ is discrete.
@@ 2. Definitions and auxiliary statements
A function $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ is called equivariant with respect to group action $\tau$ on $(\mathbb{C}^{n},0)$ if there is a character $\lambda$ of $G$ such that $f(\tau(g)\circ x)=\lambda(g)f(x)\;\forall\;x\;\forall\;g\;\in\;G$. $\lambda$ is called the weight of $f$. Invariant functions are equivariant with weight 1, of course.
@@
If $f$ is invariant with respect to an action of a group $G$, the Jacobian ideal is invariant and therefore $G$ also acts on $Q_{f}$ by changes of coordinates: $g\circ h(x)=h(g^{-1}(x))$. The representation of $G$ on $Q_{f}$ is called the equvariant Milnor number of $f$ and denoted $\mu_{G}(f)$. $\nu(f)$ is the multiplicity of the trivial representation in $\mu_{G}(f)$, i.e. the dimension of the space of invariant elements of $Q_{f}$. This subspace of invariant elements is denoted $Q^{G}_{f}$.
@@
The set $C$ of critical points of any invariant deformation $f_{\lambda}$ is $G$-invariant. The action of $G$ on $(\mathbb{C}^{n},0)$ induces a representation $\theta$ of $G$ on the space of $\mathbb{C}$-valued functions on $C$ ($\theta(g)\circ h(x)=h(g^{-1}(x))$.
Theorem 2.2.
@@
Consider the action $\tau:g\circ(x_{1},\dots,x_{n})\longrightarrow\\ \longrightarrow(x_{1},\dots,x_{k},\chi_{k+1}(g)x_{k+1},\dots,\chi_{n}(g)x_{n}),\chi_{k+1},\dots,\chi_{n}\neq 1$ and an invariant function $f$ with a zero 2jet2-jet and an isolated critical point at the origin.
@@ 3. Proof of main theorem
For a group action $\tau:g\circ(x_{1},\dots,x_{n})\longrightarrow(\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n})$ we will denote $\tau_{\mathbb{R}}$ the action on $\mathbb{C}^{2n}$ defined as $g\circ(x_{1},\dots,x_{n},y_{1},\dots,y_{n})\longrightarrow\\ (\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n},\bar{\chi}_{1}(g)y_{1},\dots,\bar{\chi}_{n}(g)y_{n})$.
$\tau_{\mathbb{R}}$ is always a real action with the invariant quadratic form $\Sigma_{i}x_{i}y_{i}$. If $\tau$ has no fixed points outside of origin then neither does $\tau_{\mathbb{R}}$.
For a function $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0) ff$$\oplus f$ is the function $f(x_{1},\dots,x_{n})+f(y_{1},\dots,y_{n})$.
It is easy to see that if $f$ has an isolated singularity at the origin then so does $f\oplus f$, and $\mu(f\oplus f)=(\mu(f))^{2}$. Furthermore, if $f$ is invariant with respect to an action $\tau$ then $f\oplus f$ is invariant with respect to the action $\tau_{\mathbb{R}}$.
@@
Now we can conclude the proof of the main theorem. As $\tau_{\mathbb{R}}$ is a real representation, we can use Roberts’ inequality for the Milnor number of $f\oplus f$:
$\mu(f\oplus f)\geq(\nu(f\oplus f)-1)l(\tau_{\mathbb{R}})+1$.
As we now know $\mu(f\oplus f)=\mu^{2}(f)$, $\nu(f\oplus f)\geq\mu(f)$ and, obviously, $l(\tau_{\mathbb{R}})=l(\tau)$. Substituting all of these into Roberts’ inequality gives us the inequality
$\mu^{2}(f)\geq(\mu(f)-1)l(\tau)+1$.
@@
4. Tightness of the bound
The next step is to produce examples showing the tightness of this bound. The trivial example is the 1-dimensional representation: if a character $\chi$ of group $G$ has order $k$ and $G$ acts on $\mathbb{C}$ via this character, the invariant function $x^{k}$ has the lowest possible Milnor number k1k-1. However, this is unsatisfactory, as we would ideally like to prove that a)the bound is tight in arbitrarily large dimensions; b)the Milnor number is bounded by the length of the shortest orbit when there are orbits of different lengths. To show this, we present the following example.
Let $n$ be an odd positive integer, $g$ a generator of a group $\mathbb{Z}_{2^{n}+1}$, $h$ a generator of $\mathbb{Z}_{3}$, $\epsilon,\sigma$ – primitive roots of unity of order $2^{n}+1,3$, respectively .
$\mathbb{Z}_{2^{n}+1}\times\mathbb{Z}_{3}$ acts on $\mathbb{C}^{n}$ in the following manner: $g\circ(x_{1},\dots,x_{n})$
$(\epsilon^{(-2)}x_{1},\epsilon^{(4)}x_{2},\dots,\epsilon^{(-2)^{n-1}}x_{n-1},\epsilon x_{n})$, $h\circ(x_{1},\dots,x_{n})=\sigma(x_{1},\dots,x_{n})$.
The polynomial $x_{1}^{2}x_{2}+x_{2}^{2}x_{3}+\dots+x_{n-1}^{2}x_{n}+x_{n}^{2}x_{1}$ is invariant and has Milnor number $2^{n}$, the length of the shortest orbit (e.g. of the point $(0,0,\dots,0,1)$) is $2^{n}+1$, however the longest orbit (e.g. of $(1,1,\dots,1)$) has length $3(2^{n}+1)$